# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 3.5 Semiconductor 2D materials
For the calculation of microscopic conductivity $\sigma_{M}(E)$ and the diffusivity $D(E)$ we use the Kubo-Greenwood formula as explained in A.1 and A.3. Defects are included directly in the Hamiltonian of the supercell. The relaxation time approximation (RTA) deals with the effect of temperature on the diffusivity through an inelastic scattering time $\tau_{i}$, which is associated to an inelastic mean free path $L_{i}$ (A.1.2). Moreover in semiconductors, it is essential to take into account the thermodynamic average due to temperature. We thus define the thermodynamic average conductivity at room temperature ($T=300$ K) as,
[…]
@@
where $N_{e}$ is the number of charge carriers with respect to the undopped system without defects. Therefore at room temperature ($T\approx 300$ K),
[…]
@@ 3.5.1 Monolayer transition metal dichalcogenides (TMDs)
There are many studies of electronic structure in TMDs (see for instance Refs. [394, 332, 333, 334, 305, 395, 335, 336, 396]), and some works on quantum transport in these new 2D materials [397, 398, 399, 400, 401, 402]. To study quantum transport, taking into account realistic scattering effects, we have developed Slater-Koster tight-binding (SK-TB) models  [294] for a series of monolayer TMDs ($MX_{2}$, where $M=$ Mo or W, and $X=$ S, Se or Te, Fig. 21(a)) in agreement with our DFT calculations performed using ABINIT software [325] (Fig. 21(d)).
[…]
@@
Figure 21: Electronic properties of monolayer MoS2. (a) Top view of the 2H TMD atomic structure $MX_{2}$: Blue and red circles indicates Metal ($M=$ Mo or W) and Chalcogen ($X=$ S, Se or Te,) atoms, respectively. The inner layer of $M$ atoms is sandwiched between two layers of $X$. (b) First Brillouin zone. (c) Total density of states (DOS) with a concentration $c$ of vacancies distributed randomly in the structure. DOS in calculated by recursion method in a 1000$\times$1000 supercell. (d) TB and DFT (ABINIT [325]) bands around the gap in MoS2 without defects. (e),(f) Thermodynamic average conductivity conductivity $\sigma_{300\,{\rm K}}$ and mobility $\mu$ at room temperature. the peak of $\mu(N_{e})$ is numerical artifact, indeed when $N_{e}\approx 0$ the equation (55) produces numerical errors because of conductivity does not go correctly to zero (tail of Lorentzian expansion). Theses calculations are performed without spin orbit coupling. $G_{0}={2e^{2}}/{h}$. Numerical method see A.3 and Eqs. (53),(55). From [294].
@@
We consider resonant defects (vacancies) randomly distributed on all atoms in an large supercell of MoS2. The main results are shown Fig. 21. Similar results for TMD serie, $MX_{2}$ ($M=$ Mo or W, $X=$ S, Se of Te), are also presented in Ref. [294]. The total density of states (DOS) in monolayer MoS2 with vacancies is shown in Fig. 21(c) for different concentration $c$ of defects with respect to the total number of atoms. The vacancies lead to the appearance of a series of peaks in the gap, which are associated to the creation of midgap states localized around the defects, whose energy and strength depends on the specific missing atoms [397, 402].
@@
Conductivity and mobility at room temperature in monolayer MoS2, calculated from equations (53),(55) and shown in Fig. 21(e,f), are in good agreement with already published results [397, 402]. Considering the TMD series [294], we found that for $n$-doped samples, $MX_{2}$ show similar mobilities, whereas for $p$-doped samples, the mobility of WTe2 is larger than for WSe2, MoTe2, MoSe2, MoS2. Numerical results show that, in general, the mobilities of TMDs are low, but they are larger for holes than for electrons, which seems to be in agreement with experimental results [403, 404, 294]. Further investigations are necessary, in particular to take account of the spin-orbit coupling, which is an essential and crucial characteristic for the applications of these semiconductor 2D materials.
@@ 3.5.2 Monolayer and multilayer black phosphorene
Figure 22: Electronic properties of monolayer, multilayers ($L=1,\cdots 5$) and bulk black phosphoerene (BP) with covalent functionalization (resonance defects) simulated by a random distribution of vacant P atoms. (a) Atomic structure of bilayer black phosphorene. A rectangular unit cell contains four P atoms per layer. (b) TB DOS of defect-free mono and multilayer phosphorene. The TB model used was proposed by Rudenko et al. [405, 406]. (c) DOS of monolayer with a concentration $c$ of vacancies. (d) Microscopic conductivity times defect concentration $c$, (bold line) $c\sigma_{M}^{xx}$ and (dashed line) $c\sigma_{M}^{yy}$ in monolayer. (e) Ratio of room temperature conductivity $\sigma_{M}^{xx}/\sigma_{M}^{yy}$ (calculated form equation (53)) versus $N_{e}$ for multilayer with vacancies concentration $c=1$%. $N_{e}$ is the density of charge carriers with respect to the neutral case. (f) Mobility at room temperature (calculated from equation (55)) along the $x$ direction versus $N_{e}$ for multilayers and bulk BP and for $c=1$%. $G_{0}={2e^{2}}/{h}$. Numerical method see A.3 and the main text. From [407].
@@
Phosphorene –monolayer and multilayer black phosphorus (BP), realized experimentally in 2014 [408, 409, 410] is a very promising 2D semiconductor due to its gap value (around $2$ eV) which decreases as the number of planes increases, reaching $\sim 0.3$ eV for bulk BP [411, 412, 413, 414, 415] (Fig. 22). This shows an increase of the transversal confinement when the number of layers decreases. Moreover, as the number of layers decreases, the electron-hole asymmetry becomes less pronounced. Indeed, for the monolayer the DOS is almost symmetric with respect to the bandgap, whereas as the number of layers increases, the DOS of the valence band becomes lower than that of the conduction band  (Fig. 22(b)). Another characteristic of monolayer and multilayer BPs is a strong in-plane structural anisotropy, which is reflected in the band anisotropy around the gap and thus in the effective masses of electrons and holes. [412, 416, 410, 417].
@@
The Phosphorene is highly sensitive to the environment and covalent or noncovalent functionalization or disorder can modify its electronic properties [405, 406, 418, 419, 420, 421, 422, 423, 424, 425, 407, 426]. Quantum diffusion calculations show that resonant and nonresonant defects act similarly on electrical conductivity [407]. In fact, even at the high defect concentrations studied, the gap states of resonant defects (Fig. 22(c)) and the gap edge states of nonresonant defects contribute very little to conductivity. As a result, the mobility gap remains nearly identical to the gap of the defect-free system. This suggests that electronic transport should be fairly well described by a semi-classical Bloch-Boltzmann-type approach, even when defects strongly modify the electronic structure. Particularly, the conductivity of states outside the gap is inversely proportional to the resonant defect concentration $c$ (Fig. 22(d)). For sufficiently small amounts of resonant defects, the calculated electron and hole mobilities are comparable to the experimental results. The strong anisotropy persists in the presence of the two types of defects we have studied and it is difficult to find very different behavior between resonant defects and nonresonant defects [407]. It is stronger in the case of monolayer BP and it is pretty close to that of the bulk BP for multilayers. We also showed that the hole conductivity is greater than that of the electrons, particularly in multilayers and bulk, and this is valid for both types of defects. This effect is related to the asymmetry in their DOS without defects.
@@ 4 Conclusion
In this review, we present some aspects on electronic structure, magnetism, and quantum transport properties of flat-band states in 2D materials and quasicrystals. We have focused on results obtained with tight-binding Hamiltonians that allow real materials to be simulated efficiently. On one hand, in perfect (defect-free) quasicrystals, complex intermetallic phases close to quasicrystals, and quasiperiodic tilings, these states characterized by very low dispersion are present in all over the spectrum, and they can even lead to a sub-diffusive regime of the charge carriers. In twisted bilayer graphene and twisted bilayer MoS2, in the absence of defects, flat-band states arising from the moiré pattern geometry are also found at low energies. In these two families of materials, the flat-band states are characterized by an extremely large spatial extent, spanning an extent over hundreds or even thousands of atoms. On the other hand, in systems (quasiperiodic tilings, graphene, transition metal dichalcogenides (TMDs), bilayer 2D materials) with resonant static defects, weakly dispersive defect states associated with a relatively large extent in real space can also exist. For instance this occurs when the defects are randomly distributed over a sublattice of the atomic structure. In both cases –whether with or without defects–, the usual approaches to study electronic transport, such as the semiclassical Bloch-Boltzmann method, are not applicable. Thus one must use methods based on the fully quantum Kubo-Greenwood formula, in real space or reciprocal space, to correctly account for all quantum effects beyond the semi-classical approximation. Our calculations show the importance of contributions beyond the Bloch-Boltzmann approximation, which are related to the non-diagonal terms of the velocity operator in the basis of Bloch states.
Regarding quasicrystals, quantum diffusion of charge carriers in 2D quasi-periodic tilings such as Penrose and octagonal tiling has numerically demonstrated the existence of sub-diffusive states, in defect-free quasiperiodic structures, even under a pure hopping Hamiltonian. At high diffusion times, these states behave like insulating states. However, it seems that in most cases, these states are not isolated and can have energies very close to super-diffusive states that behave like metallic states. It is now crucial to gain a deeper understanding of the distribution of these sub-diffusive states within the spectrum and to study in greater detail the effect of static defects and temperature on diffusion properties.
@@
In conclusion, the present review contributes to the understanding of the electronic properties of localized states (flat bands) due to the combined effect of quantum interference and geometrical properties (here moiré pattern, quasiperiodicity, or bipartite lattice). The physics of flat bands is currently a major topic in condensed matter, either for the field of topological insulators or for remarkable electronics (correlation effect, superconductivity, …). It contributes to highlight the peculiarity of the localization in these systems: a very low energy dispersion, associated with confined electronic states with long lifetimes, i.e., a small velocity, and a spatial extent of these states over a very large number of atoms (several hundred or even several thousand). This type of state leads to a particular transport regime of non-interacting charge carrier conduction which we call “small velocity regime”.
@@ Acknowledgment
The authors would like to extend their warmest thanks to our collaborating colleagues, with whom many discussions and scientific projects were carried out, Claire Berger, Duc Nguyen-Manh, Takeo Fujiwara, Pierre Mallet, Jean-Yves Veuillen, Stephan Roche, Vincent T. Renard, Claude Chapelier, Loïc Huder, Florie Mesple and T. Le Quang, Thi Thu Phùng, Robert Peters. We would also like to thank C. Oguey, A. Szállás, and A. Jagannathan, who provided us with the structures of the quasiperiodic tiling approximants studied. Most of the calculations were carried out at the Centre de Calculs (CDC), CY Cergy Paris Université, and using HPC resources from GENCI-IDRIS (Grant No. 910784). We thank Y. Costes and B. Mary, CDC, for computing assistance. GTL, AH, LM, and DM were supported by the ANR projects J2D (ANR-15-CE24-0017) and FlatMoi (ANR-21-CE30-0029). GTL, SV, JV, AM, AH, and JJK thank CY Advanced Studies, CY Cergy Paris University, for financial support.
Author contributions
GTL, JPJ, and DM developed the concept of SVR. GTL and DM studied quantum transport in 2D quasiperiodic tilings. The electronic structure of TBGs was initially studied by GTL, LM, and DM and then in collaboration with OFN and AM. Magnetism of TBGs has been studied by JV, AH, AM, and GTL. AM, JJK, DM, and GTL worked on the electronic properties of the Bernal graphene bilayers. GJ, JJK, and DM worked on the transport properties of carbon nanotubes. SV, AH, AM, and GTL led the study of twisted bilayer MoS2 and transport in monolayer TMDs. JJK, KC, GJ, DM, and GTL have studied transport in mono and mulilayer phosphorene. GTL coordinated the project with input from all authors.

### v2

@@ 3.5 Semiconductor 2D materials
For the calculation of the microscopic conductivity $\sigma_{M}(E)$ and the diffusivity $D(E)$ we use the Kubo-Greenwood formula as explained in A.1 and A.3. Defects are included directly in the Hamiltonian of the supercell. The relaxation time approximation (RTA) deals with the effect of temperature on the diffusivity through an inelastic scattering time $\tau_{i}$, which is associated to an inelastic mean free path $L_{i}$ (A.1.2). Moreover, in semiconductors, it is essential to take into account the thermodynamic average due to temperature. We thus define the thermodynamic average conductivity at room temperature ($T=300$ K) as,
$\displaystyle\sigma(\mu_{C},T\approx 300{\rm K})=\int_{-\infty}^{+\infty}{d}E\,\sigma_{M}(E)\left(-\frac{\partial f}{\partial E}\right),$ (53)
@@
where $N_{e}$ is the number of charge carriers with respect to the undoped system without defects. Therefore at room temperature ($T\approx 300$ K),
$\displaystyle\mu(\mu_{C},T\approx 300{\rm K})=\frac{1}{e|N_{e}|}\int_{-\infty}^{+\infty}{d}E\,\sigma_{M}(E)\left(-\frac{\partial f}{\partial E}\right).$ (55)
@@ 3.5.1 Monolayer transition metal dichalcogenides (TMDs)
There are many studies of electronic structure in TMDs (see for instance Refs. Huisman71 , Cappelluti13 , Rostami13 , Zahid13 , Roldan14b , Roldan14 , Ridolfi15 , SilvaGuillen16 , Angeli21 ), and some works on quantum transport in these new 2D materials Yuan14 , Stauber15 , Krstajic16 , Garcia17 , Chu19 , Gali20 . To study quantum transport, taking into account realistic scattering effects, we have developed Slater-Koster tight-binding (SK-TB) models  Venky_these for a series of monolayer TMDs ($MX_{2}$, where $M=$ Mo or W, and $X=$ S, Se or Te, Fig. 21(a)) in agreement with our DFT calculations performed using the ABINIT software Gonze16 (Fig. 21(d)).
@@
Figure 21: Electronic properties of monolayer MoS2. (a) Top view of the 2H TMD atomic structure $MX_{2}$: Blue and red circles indicate Metal ($M=$ Mo or W) and Chalcogen ($X=$ S, Se or Te,) atoms, respectively. The inner layer of $M$ atoms is sandwiched between two layers of $X$. (b) First Brillouin zone. (c) Total density of states (DOS) with a concentration $c$ of vacancies distributed randomly in the structure. The DOS is calculated by the recursion method in a $\times1000$ supercell. (d) TB and DFT (ABINIT Gonze16 ) bands around the gap in MoS2 without defects. (e),(f) Thermodynamic average conductivity conductivity $\sigma_{300\,{\rm K}}$ and mobility $\mu$ at room temperature. The peak of $\mu(N_{e})$ is a numerical artifact, indeed when $N_{e}\approx 0$ the equation (55) produces numerical errors because of conductivity does not go correctly to zero (tail of Lorentzian expansion). These calculations are performed without spin orbit coupling. $G_{0}={2e^{2}}/{h}$. For the numerical method see A.3 and Eqs. (53), (55). From Venky_these .
For a good description of the electronic bands around the gap, we include the nature of the orbitals involved, $d$ orbitals of metal ($M$) and $p$ orbitals of chalcogen ($X$). Indeed, whereas the states around the gap at the Fermi energy $E_{F}$ are mainly 4dd (5dd) states of Mo (W), a Hamiltonian restricted to those orbitals is not sufficient to recover states at $E_{F}$, since the ligand field ($X$ atoms) splits the $d$ levels of the transition metal atoms, and thus creates a direct gap at the K point Huisman71 . Therefore, all TB models proposed in the literature include at least $p$ $X$ orbitals Cappelluti13 , Rostami13 , Zahid13 , Ridolfi15 , SilvaGuillen16 . Roughly speaking the valence band has mainly $d_{0}=4d_{z^{2}}$ ($5d_{z^{2}}$) Mo (W) character, whereas the conduction band has $d_{0}$ character mixed with $d_{2}=4d_{x^{2}-y^{2}},\,4d_{xy}$ ($5d_{x^{2}-y^{2}},\,5d_{xy}$) Mo (W) character near the gap, and $d_{1}=4d_{xz},\,4d_{yz}$ ($5d_{xz},\,5d_{yz}$) Mo (W) character for higher energies Cappelluti13 . Furthermore, it seems that $3p$ ($4p$) ($5p$) S (Se) (Te) orbitals, which have lower on-site energies, act as a perturbation of the $d$ $M$ bands. For this reason, several SK-TB models Cappelluti13 , Ridolfi15 , SilvaGuillen16 fit rather well to the DFT band structure, while they propose very different sets of parameters (on-site energies and SK parameters). Our SK-TB models Venky_these , Venky20 include 11 orbitals per unit cell: 5 $d$ $M$ orbitals, $d_{0}=4d_{z^{2}}$ ($5d_{z^{2}}$), $d_{1}=4d_{xz},\,4d_{yz}$ ($5d_{xz},\,5d_{yz}$), $d_{2}=4d_{x^{2}-y^{2}},\,4d_{xy}$ ($5d_{x^{2}-y^{2}},\,5d_{xy}$) of 1 Mo (W) atom), and 6 $p$ $X$ orbitals ($3p_{x}$ ($4p_{x}$) ($5p_{x}$), $3p_{y}$ ($4p_{y}$) ($5p_{y}$) and $3p_{z}$ ($4p_{z}$) ($5p_{z}$) of 2 S (Se) (Te) atoms. Note that our SK-TB model for monolayer MoS2 is close to the model proposed in Ref. Ridolfi15 .
@@
We consider resonant defects (vacancies) randomly distributed on all atoms in a large supercell of MoS2. The main results are shown Fig. 21. Similar results for the TMD series, $MX_{2}$ ($M=$ Mo or W, $X=$ S, Se of Te), are also presented in Ref. Venky_these . The total density of states (DOS) in monolayer MoS2 with vacancies is shown in Fig. 21(c) for different concentration $c$ of defects with respect to the total number of atoms. The vacancies lead to the appearance of a series of peaks in the gap, which are associated to the creation of midgap states localized around the defects, whose energy and strength depends on the specific missing atoms Yuan14 , Gali20 .
@@
Conductivity and mobility at room temperature in monolayer MoS2, calculated from equations (53), (55) and shown in Fig. 21(e,f), are in good agreement with already published results Yuan14 , Gali20 . Considering the TMD series Venky_these , we found that for $n$-doped samples, $MX_{2}$ show similar mobilities, whereas for $p$-doped samples, the mobility of WTe2 is larger than for WSe2, MoTe2, MoSe2, or MoS2. Numerical results show that, in general, the mobilities of TMDs are low, but they are larger for holes than for electrons, which seems to be in agreement with experimental results Zhu14 , Chu14 , Venky_these . Further investigations are necessary, in particular to take account of the spin-orbit coupling, which is an essential and crucial characteristic for the applications of these semiconductor 2D materials.
@@ 3.5.2 Monolayer and multilayer black phosphorene
Figure 22: Electronic properties of monolayer, multilayers ($L=1,\cdots 5$) and bulk black phosphoerene (BP) with covalent functionalization (resonant defects) simulated by a random distribution of vacant P atoms. (a) Atomic structure of bilayer black phosphorene. A rectangular unit cell contains four P atoms per layer. (b) TB DOS of defect-free mono and multilayer phosphorene. The TB model used was proposed by Rudenko et al. Rudenko2014 , Rudenko2015 . (c) DOS of a monolayer with a concentration $c$ of vacancies. (d) Microscopic conductivity times defect concentration $c$, (bold line) $c\sigma_{M}^{xx}$ and (dashed line) $c\sigma_{M}^{yy}$ in a monolayer. (e) Ratio of room temperature conductivity $\sigma_{M}^{xx}/\sigma_{M}^{yy}$ (calculated form equation (53)) versus $N_{e}$ for multilayers with vacancies concentration c=1c=1%. $N_{e}$ is the density of charge carriers with respect to the neutral case. (f) Mobility at room temperature (calculated from equation (55)) along the $x$ direction versus $N_{e}$ for multilayers and bulk BP and for c=1c=1%. $G_{0}={2e^{2}}/{h}$. For the numerical method see A.3 and the main text. From Jouda24 .
@@
Phosphorene –monolayer and multilayer black phosphorus (BP), realized experimentally in 2014 Li2014 , Liu2014 , Xia2014 is a very promising 2D semiconductor due to its gap value (around $2$ eV) which decreases as the number of planes increases, reaching $\sim 0.3$ eV for bulk BP Tran2014 , Qiao2014 , cai2014layer , li2017direct , Gaufres19 (Fig. 22). This shows an increase of the transversal confinement when the number of layers decreases. Moreover, as the number of layers decreases, the electron-hole asymmetry becomes less pronounced. Indeed, for the monolayer the DOS is almost symmetric with respect to the band gap, whereas as the number of layers increases, the DOS of the valence band becomes lower than that of the conduction band  (Fig. 22(b)). Another characteristic of monolayer and multilayer BPs is a strong in-plane structural anisotropy, which is reflected in the band anisotropy around the gap and thus in the effective masses of electrons and holes Qiao2014 , Carvalho2016 , Xia2014 , Akhtar2017 .
@@
Phosphorene is highly sensitive to the environment and covalent or noncovalent functionalization or disorder can modify its electronic properties Rudenko2014 , Rudenko2015 , Paez16 , Mitrovic21 , ryder2016 , tang2017fluorinated , jellett2020prospects , van2018covalent , yan2019hydroxyl , Guzman23 , Jouda24 , Oubram25 . Quantum diffusion calculations show that resonant and nonresonant defects act similarly on the electrical conductivity Jouda24 . In fact, even at the high defect concentrations studied, the gap states of resonant defects (Fig. 22(c)) and the gap edge states of nonresonant defects contribute very little to the conductivity. As a result, the mobility gap remains nearly identical to the gap of the defect-free system. This suggests that electronic transport should be fairly well described by a semi-classical Bloch-Boltzmann-type approach, even when defects strongly modify the electronic structure. Particularly, the conductivity of states outside the gap is inversely proportional to the resonant defect concentration $c$ (Fig. 22(d)). For sufficiently small amounts of resonant defects, the calculated electron and hole mobilities are comparable to the experimental results. The strong anisotropy persists in the presence of the two types of defects we have studied and it is difficult to find very different behavior between resonant defects and nonresonant defects Jouda24 . It is stronger in the case of monolayer BP and it is pretty close to that of bulk BP or multilayers. We also showed that the hole conductivity is greater than that of the electrons, particularly in multilayers and the bulk, and this is valid for both types of defects. This effect is related to the asymmetry in their DOS without defects.
@@ 4 Conclusion
In this review, we present some aspects on electronic structure, magnetism, and quantum transport properties of flat-band states in 2D materials and quasicrystals. We have focused on results obtained with tight-binding Hamiltonians that allow real materials to be simulated efficiently. On one hand, in perfect (defect-free) quasicrystals, complex intermetallic phases close to quasicrystals, and quasiperiodic tilings, these states characterized by very low dispersion are present throughout the spectrum, and they can even lead to a sub-diffusive regime of the charge carriers. In twisted bilayer graphene and twisted bilayer MoS2, in the absence of defects, flat-band states arising from the moiré pattern geometry are also found at low energies. In these two families of materials, the flat-band states are characterized by an extremely large spatial extent, spanning an extent over hundreds or even thousands of atoms. On the other hand, in systems (quasiperiodic tilings, graphene, transition metal dichalcogenides (TMDs), multilayer 2D materials) with resonant static defects, weakly dispersive defect states associated with a relatively large extent in real space can also exist. For instance this occurs when the defects are randomly distributed over a sublattice of the atomic structure. In both cases –whether with or without defects–, the usual approaches to study electronic transport, such as the semiclassical Bloch-Boltzmann method, are not applicable. Thus one must use methods based on the fully quantum Kubo-Greenwood formula, in real space or reciprocal space, to correctly account for all quantum effects beyond the semi-classical approximation. Our calculations show the importance of contributions beyond the Bloch-Boltzmann approximation, which are related to the non-diagonal terms of the velocity operator in the basis of Bloch states.
Regarding quasicrystals, quantum diffusion of charge carriers in 2D quasi-periodic tilings such as Penrose and octagonal tilings has numerically demonstrated the existence of sub-diffusive states, in defect-free quasiperiodic structures, even under a pure hopping Hamiltonian. At high diffusion times, these states behave like insulating states. However, it seems that in most cases, these states are not isolated and can have energies very close to super-diffusive states that behave like metallic states. It is now crucial to gain a deeper understanding of the distribution of these sub-diffusive states within the spectrum and to study in greater detail the effect of static defects and temperature on diffusion properties.
@@
In conclusion, the present review contributes to the understanding of the electronic properties of localized states (flat bands) due to the combined effect of quantum interference and geometrical properties (here moiré pattern, quasiperiodicity, or bipartite lattice). The physics of flat bands is currently a major topic in condensed matter, either for the field of topological insulators or for remarkable electronics (correlation effect, superconductivity, …). It contributes to highlight the peculiarity of the localization in these systems: a very low dispersion, associated with confined electronic states with long lifetimes, i.e., a small velocity, and a spatial extent of these states over a very large number of atoms (several hundred or even several thousand). This type of state leads to a particular transport regime of non-interacting charge carrier conduction which we call “small velocity regime”.
@@ Acknowledgment
The authors would like to extend their warmest thanks to our collaborating colleagues, with whom many discussions and scientific projects were carried out, Claire Berger, Duc Nguyen-Manh, Takeo Fujiwara, Pierre Mallet, Jean-Yves Veuillen, Stephan Roche, Vincent T. Renard, Claude Chapelier, Loïc Huder, Florie Mesple and T. Le Quang, Thị Thu Phùng, Robert Peters. We would also like to thank C. Oguey, A. Szállás, and A. Jagannathan, who provided us with the structures of the quasiperiodic tiling approximants studied. Most of the calculations were carried out at the Centre de Calculs (CDC), CY Cergy Paris Université, and using HPC resources from GENCI-IDRIS (Grant No. 910784). We thank Yann Costes and Baptiste Mary, CDC, for computing assistance. JJK was supported by the Tunisian French Cooperation Project (CMCU 15G1306). GTL, AH, LM, and DM were supported by the ANR projects J2D (ANR-15-CE24-0017) and FlatMoi (ANR-21-CE30-0029). GTL, SV, JV, AM, AH, and JJK thank CY Advanced Studies, CY Cergy Paris University, for financial support.